OpenStudy (anonymous):

which of the following are equal to the expression sqrt9/sqrt81 a. 3/9 b.3 c.2/9 d.1/3 ?? help please ?

OpenStudy (anonymous):

d,a

OpenStudy (anonymous):

\[\huge \frac{\sqrt9}{\sqrt{81}}=\frac{\sqrt{3^2}}{\sqrt{9^2}}=\frac39=\frac13\]